OpenStudy (anonymous):

Write your result in standard form: 3i(4-5i) + 2i(4i-√-25)

OpenStudy (asnaseer):

which part of this are you stuck on?

OpenStudy (anonymous):

I just don't understand where to start.

OpenStudy (asnaseer):

ok, do you know how to expand each of the bracketed expressions?

OpenStudy (anonymous):

I don't :(

OpenStudy (asnaseer):

ok, lets start from the beginning then. if we have: a(b+c) this means "a" multiplied by the sum of "b" and "c". this gives you "a" lots of "b" and "c", so we can expand this as follows: a(b+c) = ab + ac do you understand this part?

OpenStudy (anonymous):

Yes, I do.

OpenStudy (asnaseer):

e.g.: 3(4+2) = 3*4 + 3*2 = 12 + 6 = 18 we can this as 4+2=6, so 3(4+2) = 3 * 6 = 18

OpenStudy (asnaseer):

ok, so lets try to expand the first expression: \[3i(4-5i) =3i*4 - 3i*5i=12i-15i^2\]now since \(i=\sqrt{-1}\) we know that \(i^2=-1\), we therefore get:\[3i(4-5i)=12i-15i^2=12i-(15*-1)=12i-(-15)=12i+15\] got that?

OpenStudy (anonymous):

yes, i do. the radicals are what confuse me.

OpenStudy (asnaseer):

ok, so lets look at those nasty radicals next...

OpenStudy (asnaseer):

in this expression:\[2i(4i-\sqrt{-25})\]the \(\sqrt{-25}\) can be written as follows:\[\sqrt{-25}=\sqrt{25*-1}=\sqrt{25}*\sqrt{-1}=\sqrt{25}*i=5i\]so we can rewrite the original as:\[2i(4i-5i)\]understand so far?

OpenStudy (anonymous):

Where did you get the -1? Or do we multiply by that so we can get rid of the negative?

OpenStudy (asnaseer):

-25 = 25 * -1

OpenStudy (asnaseer):

any positive number times -1 will give you the negative of that number

OpenStudy (anonymous):

Right, I understand that now.

OpenStudy (asnaseer):

\[2i(4i-\sqrt{-25})=2i(4i-5i)=2i(-i)=2i * -i = -2i^2 = -2 * i^2\]\[= -2 * (-1) = 2\]

OpenStudy (asnaseer):

does that make sense?

OpenStudy (anonymous):

yes.

OpenStudy (asnaseer):

ok, so finally we put it all together to get:\[12i+15+2=12i+17\]the standard form for a complex number is usually written as:\[a+bi\]where \(i=\sqrt{-1}\) so we can just rearrange the final answer to get:\[17+12i\]

OpenStudy (anonymous):

That was a tough one! I'm starting to get the hang of this radical stuff. Thank you so much.

OpenStudy (asnaseer):

yw - I'm glad I could help you understand.

OpenStudy (anonymous):

I'm terrible at this math stuff. I'm glad there are people like you that can help break it down.

OpenStudy (asnaseer):

have faith in your abilities and you will achieve things you never thought you could. :-)

OpenStudy (anonymous):

I will surely remember that :)

OpenStudy (asnaseer):

take care...

OpenStudy (anonymous):

you too! Thank you once again.